OpenStudy (anonymous):

solve the equation. check the solution.-4/x+1=-1/x+5

OpenStudy (vickyisthesmartone123):

So do we solve for X?

OpenStudy (boldjon):

x=-3/4

OpenStudy (serenity74):

Is it like -4/(x+1)= -1/(x+5) cause that matters.

OpenStudy (vickyisthesmartone123):

@boldjon EXACTLY :)

OpenStudy (vickyisthesmartone123):

Do you have answer choices @chubb108 ?

OpenStudy (anonymous):